OpenStudy (anonymous):

the initial velocity of a particle is u ta t=0, and the acceleration is given by f=at.which of the following relation is valid a)v=u+at2 b)v=u+at2/2 c)v=u=at d)v=u

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

yes it is but how...??????????

OpenStudy (anonymous):

because the initial velocity is o and the the velocity is only depends on multyple of acc and time...in b there 2/2=1; hence v=u+at;u=0

OpenStudy (anonymous):

listen here the b) v=u+at^2/2

OpenStudy (anonymous):

sorry,, i couldnt see that..

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

\[acceleration=dv/dt=(at)=given\] \[\implies dv=(at)dt\] \[\int\limits_{u}^{v(final)}dv=\int\limits_{0}^{t(final)}(at)dt\] \[v(final)=u+at ^{2}/2\]